1. Consider the production function f(x1, x2) = 4x 11/2 x 21/3 . What is the returns to scale? Show your work.

2. What is the TRS for the above production function?

3. What is the optimal level of output that maximizes profit given the output and input prices respectively as p, w1, w2?
Answer #1

© The given production function is - Q = f(x1, x2) = 4 x 22:43 Now, f (+X12+x2) = 4 (tx, 12(tx2)? t> 0 4+3+5 27,!2X2!3 = 481xy + X20X2] = 0 9 = 1[*312, dX2 112 © Talithal Reese of a tribuchon (TRS)= ab Tally preferentialing the production function v